# Zero-shot Generalizable Incremental Learning for Vision-Language Object Detection

### v1

@@ Abstract
This paper presents Incremental Vision-Language Object Detection (IVLOD), a novel learning task designed to incrementally adapt pre-trained Vision-Language Object Detection Models (VLODMs) to various specialized domains, while simultaneously preserving their zero-shot generalization capabilities for the generalized domain. To address this new challenge, we present the Zero-interference Reparameterizable Adaptation (ZiRa), a novel method that introduces Zero-interference Loss and reparameterization techniques to tackle IVLOD without incurring additional inference costs or a significant increase in memory usage. Comprehensive experiments on COCO and ODinW-13 datasets demonstrate that ZiRa effectively safeguards the zero-shot generalization ability of VLODMs while continuously adapting to new tasks. Specifically, after training on ODinW-13 datasets, ZiRa exhibits superior performance compared to CL-DETR and iDETR, boosting zero-shot generalizability by substantial 13.91 and 8.71 AP, respectively.
@@ 1 Introduction
Object detection has achieved remarkable strides in recent years [2, 39]. However, most object detection models are typically trained to predict a predefined, closed set of categories [29, 2], constraining their prospect in real-world applications. Recent research endeavors [11, 24, 38, 17, 20] focus on developing intelligent systems capable of detecting objects specified by natural language inputs, giving rise to the field of Vision-Language Object Detection (VLOD), wherein the models developed are named Vision-Language Object Detection Models (VLODMs) accordingly. The inputted language encompasses a broader array of visual concepts than a predefined concept set, leading to outstanding zero-shot generalizability.
Figure 1: IVLOD aims to enhance VLODMs’ performance across specialized domains via incremental learning, while also preserving their zero-shot generalization capability, enabling them to handle both known and unknown objects simultaneously and effectively.
@@
A straightforward idea is to adapt different individual VLODMs to different downstream tasks. However, a general agent often needs to simultaneously recognize objects from diverse tasks, each learned at distinct times. Take unmanned supermarkets as an example: pre-setting all potential future categories is not practical, and there is an urgent necessity for a robot to simultaneously recognize all categories of goods learned over time. Employing multiple models for each task falls short of this demand, making it crucial to incrementally learn with a single adaptable model. Moreover, in a dynamically changing environment, equipping VLODMs with the ability of incremental learning is paramount [18, 28], considering that downstream tasks typically arrive in a streaming manner and incremental learning paradigm better aligns with the learning and cognition processes of a general intelligent agent.
@@
To effectively address IVLOD’s challenges while improving memory efficiency, we introduce a novel approach named Zero-interference Reparameterizable Adaptation (ZiRa), comprising the following two components.
Reparameterizable Dual Branch (RDB). A fundamental difference between continual learning in the general setting and IVLOD is that IVLOD starts from a well-pre-trained model. Existing research convincingly demonstrates that effective adaptation to downstream tasks can be achieved by tuning only a subset of additional parameters [13, 14]. ZiRa seizes upon this insight by retaining the parameters of the original model and introducing a parallel dual branch structure RDB designed for efficient tuning on downstream tasks. This RDB structure serves a dual purpose—it achieves the branch labor division for downstream continual learning, protecting the learned knowledge from being excessively overwritten, and more crucially, it lays the structural foundation for the second new component.
Zero-interference Loss (ZiL). Based on the RDB’s dual-branch structure, ZiRa further incorporates a novel pivotal element named ZiL. The purpose of ZiL is to reduce the interference of new knowledge on learned knowledge. To achieve this, ZiL simultaneously penalizes the output norms of both the entire RDB and a high-learning-rate branch within the RDB, which guides the RDB to learn in a direction that protects both knowledge learned from pre-training and downstream tasks. As a result, ZiL effectively addresses the two challenges of IVLOD at the same time. Crucially, ZiRa accomplishes this without necessitating the duplication of the entire model or any exemplars, thereby ensuring efficient utilization of memory.
Compared to existing methods [31, 21], ZiRa leverages the advantages offered by VLODMs more fully through a principled design (i.e., the ZiL based on RDB), offering a superior and more effective solution without the need for excessive additional memory and inference cost. The contributions of this paper can be summarized as follows:
@@
  • We conduct comprehensive experiments on COCO and ODinW-13 datasets, showcasing the effectiveness of our method. For instance, ZiRa outperforms CL-DETR [21] and iDETR [9] by a significant margin, making a 13.91 and 8.71 improvement in zero-shot AP on COCO, respectively, while offering competitive results on downstream tasks incremental learning.
@@ 2 Related Work
Vision-Language Object Detection. Vision-Language Object Detection Models (VLODMs) combine natural language descriptions with object detection in images. Early research in this area primarily focused on grounding textual phrases to image regions, such as grounding referring expressions [25, 35, 3] and visual question answering [1, 23, 37]. More recent studies have developed unified and scalable approaches for simultaneous object detection and grounding tasks. For instance, GLIP [17] unifies phrase grounding and object detection tasks by treating object detection as context-free phrase grounding and treating phrase grounding as a contextualized object detection task. This approach enables the integration of multi-layered textual information into the detection model. Grounding DINO [20], based on the DINO [39], employs a DETR-like [2] architecture that uses a Transformer as the detector, which allows for unified processing of image and language data and offers the outstanding capacity to utilize large-scale datasets. Based on Grounding DINO, our work further addresses the newly introduced IVLOD tasks.
Incremental Learning. Incremental learning, also known as continual learning or lifelong learning, tackles the challenge of learning new tasks without forgetting previous ones. Incremental learning has given rise to two primary strategies, memory replay, and knowledge distillation, to tackle its inherent challenges [18, 28, 4, 34, 33]. For example, Learning without Forgetting (LwF) [18] employs knowledge distillation to maintain performance on old tasks while learning new ones. Another notable approach, iCaRL [28], combines exemplar memory and a nearest-mean-of-exemplars classification method for incremental learning.
@@
Incremental Object Detection. Incremental object detection, a specialized subfield within incremental learning, focuses on improving the adaptability of object detection models to new object categories while maintaining performance on previously learned categories [31, 15, 27, 36, 10, 9, 21]. ILOD [31] was the first to introduce incremental object detection. More recently, CL-DETR [21] effectively combined knowledge distillation and exemplar memory techniques to enhance incremental object detection using Transformer-based models. Our work distinguishes itself by performing incremental learning on VLODMs, which are more favorable for open-world problems. Besides the need to continually adapt across multiple specialized tasks, it is also important to preserve their zero-shot generalization ability. To the best of our knowledge, our work is the first to pose and tackle IVLOD tasks.
@@
Reparameterization. Reparameterization techniques optimize model inference speed by transforming the training-time architecture into an equivalent, simplified form for inference [6, 7, 8]. RepVGG [7] exemplified this approach by employing $3\times 3$ and $1\times 1$ convolutional branches for training, which are streamlined to $3\times 3$ convolutional branches and ReLU activations during inference. RepMLPNet [8] introduced “Locality Injection” to merge trained parameters of parallel convolutional kernels into fully connected layers, enhancing MLP-based model performance. ZiRa not only reduces inference costs and model size through reparameterization but also synergizes the reparameterization technique with the ZiL to effectively mitigate forgetting in downstream tasks without significant memory overhead, which stands as a novel endeavor in the field of incremental learning.
@@ 3 Methodology
Figure 2: Our framework, features two Reparameterizable Dual Branch with Zero-inference Loss on both the vision and language sides.
@@ 3.2 Reparameterizable Dual Branch
Figure 3: Two sub-branches (HLRB, LLRB) inside the Reparameterizable Dual Branch (RDB). The RDB and the pre-trained branch are parallel.
Figure 3: Two sub-branches (HLRB, LLRB) inside the Reparameterizable Dual Branch (RDB). The RDB and the pre-trained branch are parallel.
@@
Single Branch for Inference. Sub-branches inside the RDB like the LLRB and HLRB can increase computational overhead. We efficiently address this by reparameterizing all sub-branches in RDB and the frozen pre-trained branch into a new one for inference. Since the frozen pre-trained branch is also a linear structure, after learning all the tasks, we can create a new branch that merges the frozen pre-trained branch and the RDB via reparameterization. During inference, we exclusively use this newly consolidated branch, but we still retain the parameters frozen pre-trained branch and the RDB for future learning.
@@ 3.3 Zero-interference Loss
When using the RDB to adapt to new tasks, if we enforce the output of the RDB to be zero, the intermediate prompts and features generated by the newly trained VLODM will remain consistent with those of the original pre-trained model. In other words, prompts and features with a zero RDB output are equivalent to those produced by the original pre-trained model. ZiRa’s fundamental design relies on constraining the RDB’s output to be zero through ZiL. Thus, if we optimize ZiL, the optimization of RDB parameters will be guided in a direction that minimizes interference with the VLODM’s zero-shot generalizability, efficiently preventing forgetting on the original pre-training domain without necessitating space to store an entire model copy for distillation or old samples for replaying.
Specifically, ZiL penalizes the norm of the whole RDB’s output (ref. Fig. 4), aiming at effectively safeguarding the zero-shot performance of VLODM. The ZiL for the RDB, $L_{rdb}$, is defined as follows:
@@
Further Discussion about ZiL. To further investigate the effectiveness of ZiL, we plot curves illustrating the evolution of the RDB’s output norm throughout the IVLOD process, as shown in Fig. 4. It is evident that training the RDB with ZiL leads to a substantial reduction in the output norm of the RDB on COCO, compared to training the RDB without ZiL. This preliminary observation strongly suggests that ZiL effectively preserves the zero-shot generalizability of VLODMs.
Examining the evolution curves in Fig. 4 more closely, we observe that in the initial stages, ZiL plays a relatively minor role in curbing the growth of the RDB’s output norm. However, as the process of accumulating new knowledge progresses, the influence of ZiL becomes more significant, leading to stronger constraint effects. This increase in ZiL’s impact leads to a dynamic balance: the interference caused by integrating new knowledge is counteracted by the interference reduction achieved by ZiL, which is reflected in the fact that as the number of tasks increases, the norm of the RDB’s output doesn’t rise unboundedly but rather stabilizes after learning a certain number of tasks.
@@ 4.1 Setup
Datasets. We conduct our experiments on the COCO [19] datasets and the “Object Detection in the Wild (ODinW)” [16] benchmark. ODinW is a more challenging benchmark designed to test model performance under real-world scenarios. It comprises numerous sub-datasets from various domains for evaluation, such as Thermal (to detect objects in heat map images) and Aquarium (to detect different marine animals). Following GLIP [17], we use ODinW-13 datasets, they are labeled as Ae (Aerial Maritime Drone), Aq (Aquarium), Co (Cottontail Rabbits), Eg (Egohands), Mu (Mushrooms), Pa (Packages), Pv (Pascal VOC), Pi (Pistols), Po (Pothole), Ra (Raccoon), Sh (Shellfish), Th (Thermal Dogs and People), Ve (Vehicles).
Training and Testing Protocols. The 13 sub-datasets of ODinW-13 are trained sequentially, one by one, and are tested after all sub-datasets have been trained. In this paper, we primarily focus on the task-agnostic setting of incremental learning, in which there is no clear task boundary among the ODinW-13 datasets and COCO. This means that the dataset ID will not be provided and the model should predict all learned classes across all datasets and domains during testing.
Metrics. As COCO is a dataset within the realm of everyday life, we use zero-shot COCO performance as a metric of the model’s zero-shot generalization ability across general domains. Additionally, we provide the average performance on the downstream ODinW-13 benchmark, with performance calculations made only after the model has been sequentially trained on all sub-datasets. We denote the zero-shot performance on COCO as “ZCOCO” and the average AP on ODinW-13 downstream tasks as “Avg”.
Implementation Details. In this paper, unless otherwise specified, we use the Grounding DINO [20] with the Swin-Tiny [22] backbone as the original VLODM in our experiments. This Grounding DINO is pre-trained on Objects365 [30], GoldG [17], and Cap4M [17]. The learning order of the ODinW-13 sub-datasets is randomly shuffled. All results are obtained after running experiments three times with different random seeds. Due to limited space, please refer to the appendix for more details on the training hyperparameters.
@@
Table 1: IVLOD results on ODinW-13 and ZCOCO. All results are based on the same VLODM. “Shots” means how many samples are used when adaptation training. “Avg” is the average AP on all ODinW-13 sub-datasets.

### v3

@@ Abstract
This paper presents Incremental Vision-Language Object Detection (IVLOD), a novel learning task designed to incrementally adapt pre-trained Vision-Language Object Detection Models (VLODMs) to various specialized domains, while simultaneously preserving their zero-shot generalization capabilities for the generalized domain. To address this new challenge, we present the Zero-interference Reparameterizable Adaptation (ZiRa), a novel method that introduces Zero-interference Loss and reparameterization techniques to tackle IVLOD without incurring a significant increase in memory usage. Comprehensive experiments on COCO and ODinW-13 datasets demonstrate that ZiRa effectively safeguards the zero-shot generalization ability of VLODMs while continuously adapting to new tasks. Specifically, after training on ODinW-13 datasets, ZiRa exhibits superior performance compared to CL-DETR and iDETR, boosting zero-shot generalizability by substantial 13.91 and 8.74 AP, respectively. Our code is available at https://github.com/JarintotionDin/ZiRaGroundingDINO.
@@ 1 Introduction
Figure 1: Incremental Vision-Language Object Detection (IVLOD) aims to enhance VLODMs’ performance across specialized domains via incremental learning, while also preserving their zero-shot generalization capability, enabling them to handle both known and unknown objects simultaneously and effectively.
Object detection has achieved remarkable strides in recent years [2, 44]. However, most object detection models are typically trained to predict a predefined, closed set of categories [32, 2], constraining their prospect in real-world applications. Recent research endeavors [12, 27, 45, 19, 22] focus on developing intelligent systems capable of detecting objects specified by natural language inputs, giving rise to the field of Vision-Language Object Detection (VLOD) (also named Open Vocabulary Object Detection), also referred to as Open Vocabulary Object Detection. In this framework, the models are termed Vision-Language Object Detection Models (VLODMs). By incorporating natural language inputs, VLODMs can recognize a much broader set of visual concepts beyond a fixed category set, resulting in exceptional zero-shot generalizability.
@@
A straightforward idea is to adapt different individual VLODMs to different downstream tasks. However, a general agent often needs to simultaneously recognize objects from diverse tasks, each learned at distinct times. Moreover, in a dynamically changing environment, equipping VLODMs with the ability of incremental learning is paramount [20, 31], considering that downstream tasks typically arrive in a streaming manner and incremental learning paradigm better aligns with the learning and cognition processes of a general intelligent agent.
@@
To effectively address IVLOD’s challenges while improving memory efficiency, we introduce a novel approach named Zero-interference Reparameterizable Adaptation (ZiRa). ZiRa retains the parameters of the original model and introduces a parallel dual branch structure named Reparameterizable Dual Branch (RDB) designed for efficient tuning on downstream tasks. The RDB structure is reparameterizable, allowing the model to adapt to downstream tasks without increasing memory usage. Furthermore, the RDB structure serves a dual purpose—it achieves the branch labor division for downstream continual learning, protecting the learned knowledge from being excessively overwritten, and more crucially, it lays the structural foundation for the key novel element of ZiRa named Zero-interference Loss (ZiL).
ZiL simultaneously penalizes the output norms of both the entire RDB and a high-learning-rate branch within the RDB, guiding the RDB to learn in a direction that protects both knowledge learned from pre-training and downstream tasks. The idea of ZiL comes from the following insight: During pre-training, the model is trained with visual-language inputs that contain significant noise, and the model has learned how to handle this noise. As a result, the pre-trained VLODMs have a certain degree of robustness to the input, and they can handle a certain range of interference without their performance being affected. ZiL ensures that the fine-tuned residual input of RDB is of small norm, thereby guaranteeing that the model’s original performance. At the same time, for adaptation, only a small adjustment to the input is needed to learn new concepts. Although ZiL constrains the input norm of the RDB, it still allows the RDB to acquire sufficient downstream knowledge. As a result, ZiL effectively addresses the two challenges of IVLOD at the same time. Crucially, ZiRa accomplishes this without necessitating the duplication of the entire model or any exemplars, thereby ensuring efficient utilization of memory.
Compared to existing methods [34, 23], ZiRa leverages the advantages offered by VLODMs more fully through a principled design (i.e., the ZiL based on RDB), offering a superior and more effective solution without the need for excessive additional memory. The contributions of this paper can be summarized as follows:
@@
  • We conduct comprehensive experiments on COCO and ODinW-13 datasets, showcasing the effectiveness of our method. For instance, ZiRa outperforms CL-DETR [23] and iDETR [10] by a significant margin, making a 13.91 and 8.74 improvement in zero-shot AP on COCO, respectively, while offering competitive results on downstream tasks incremental learning.
@@ 2 Related Work
Vision-Language (or Open Vocabulary) Object Detection. Vision-Language (or Open Vocabulary) Object Detection Models combine natural language descriptions with object detection in images. Early research in this area primarily focused on grounding textual phrases to image regions, such as grounding referring expressions [28, 39, 4] and visual question answering [1, 25, 42]. More recent studies have developed unified and scalable approaches for simultaneous object detection and grounding tasks. For instance, GLIP [19] unifies phrase grounding and object detection tasks by treating object detection as context-free phrase grounding and treating phrase grounding as a contextualized object detection task. This approach enables the integration of multi-layered textual information into the detection model. Grounding DINO [22], based on the DINO [44], employs a DETR-like [2] architecture that uses a Transformer as the detector, which allows for unified processing of image and language data and offers the outstanding capacity to utilize large-scale datasets. Based on Grounding DINO, our work further addresses the newly introduced IVLOD tasks.
@@
Incremental Learning and Incremental Object Detection. Incremental learning, also known as continual learning or lifelong learning, tackles the challenge of learning new tasks without forgetting previous ones. Incremental learning has given rise to two primary strategies, memory replay, and knowledge distillation, to tackle its inherent challenges [20, 31, 3, 5, 38, 37, 40]. For example, Learning without Forgetting (LwF) [20] employs knowledge distillation to maintain performance on old tasks while learning new ones. Another notable approach, iCaRL [31], combines exemplar memory and a nearest-mean-of-exemplars classification method for incremental learning. Incremental object detection, a specialized subfield within incremental learning, focuses on improving the adaptability of object detection models to new object categories while maintaining performance on previously learned categories [34, 17, 30, 41, 11, 10, 23]. ILOD [34] was the first to introduce incremental object detection. More recently, CL-DETR [23] effectively combined knowledge distillation and exemplar memory techniques to enhance incremental object detection using Transformer-based models. Our work distinguishes itself by performing incremental learning on VLODMs, which are more favorable for open-world problems. Besides the need to continually adapt across multiple specialized tasks, it is also important to preserve their zero-shot generalization ability. To the best of our knowledge, our work is the first to pose and tackle IVLOD tasks.
Open-World Object Detection. Open-World Object Detection (OWOD) [16, 13, 26] also targets the detection of both seen and unseen objects, but it involves simultaneous learning of new objects and detecting unknown objects during the incremental learning process. This means that as new tasks are introduced, the model is trained to recognize new objects while also developing its ability to detect unknown objects. In contrast, Incremental Vision-Language Object Detection (IVLOD) pre-trains models to detect unknown objects before the incremental learning process begins. The focus of IVLOD during incremental learning is on preserving the model’s ability to detect these unknown objects while adapting to new tasks. Given the superior zero-shot generalization capabilities of current vision-language models, we believe that starting with a pre-trained vision-language model for IVLOD is a more effective approach compared to the traditional OWOD paradigm.
@@
Reparameterization. Reparameterization techniques optimize model inference speed by transforming the training-time architecture into an equivalent, simplified form for inference [8, 9, 7]. RepVGG [9] exemplified this approach by employing $3\times 3$ and $1\times 1$ convolutional branches for training, which are streamlined to $3\times 3$ convolutional branches and ReLU activations during inference. RepMLPNet [7] introduced “Locality Injection" to merge trained parameters of parallel convolutional kernels into fully connected layers, enhancing MLP-based model performance. Our work builds upon the Reparameterizable Dual Branch (RDB) concept initially proposed by [43]. The key innovations introduced in our approach include the use of differentiated learning rates for the High Learning Rate Branch and Low Learning Rate Branch, and the implementation of Zero-interference Loss (ZiL). These enhancements allow for a more effective balance between learning new tasks and preserving previously acquired knowledge, distinguishing our method from prior work.
@@ 3.1 Overview
Figure 2: Our framework, features two Reparameterizable Dual Branch with Zero-interference Loss on both the vision and language sides.
@@ 3.2 Reparameterizable Dual Branch
Single Branch for Inference. Sub-branches inside the RDB like the LLRB and HLRB can increase computational overhead. We efficiently address this by reparameterizing all sub-branches in RDB and the frozen pre-trained branch into a new one for inference. Since the frozen pre-trained branch is also a linear structure, after learning all the tasks, we can create a new branch that merges the frozen pre-trained branch and the RDB via reparameterization. During inference, we exclusively use this newly consolidated branch, but we still retain the parameters frozen pre-trained branch and the RDB for future learning. This approach significantly reduces the computational cost during inference, as only one branch is required for prediction
@@ 3.3 Zero-interference Loss
Specifically, ZiL penalizes the norm of the whole RDB’s output (ref. Fig. 3), aiming at safeguarding the zero-shot performance of VLODM. The ZiL for the RDB, $L_{rdb}$, is defined as follows:
@@
Why ZiL works. The effectiveness of ZiL can be explained as follows. First, the VLODM’s input (What we fine-tuned in this paper) is robust to small-norm random perturbations, which have little impact on the model’s performance. To verify this, we add Gaussian noise to the input of the pre-trained VLODM’s detector and observe the model’s performance. As shown in Fig. 5, the model’s performance is not significantly affected by the addition of small-norm random noise to the input.
Figure 4: The performance of the pre-trained VLODM with different levels of Gaussian noise added to the input of VLODM’s detector.
Second, the ZiL ensures that the fine-tuned input of the VLODM has a small norm additional term. To further investigate the effectiveness of ZiL, we plot curves illustrating the evolution of the RDB’s output norm throughout the IVLOD process, as shown in Fig. 5. It is evident that training the RDB with ZiL leads to a substantial reduction in the output norm of the RDB on COCO, compared to training the RDB without ZiL. This preliminary observation strongly suggests that ZiL effectively preserves the zero-shot generalizability of VLODMs. Examining the evolution curves in Fig. 5 more closely, we observe that in the initial stages, ZiL plays a relatively minor role in curbing the growth of the RDB’s output norm. However, as the process of accumulating new knowledge progresses, the influence of ZiL becomes more significant, leading to stronger constraint effects. This increase in ZiL’s impact leads to a dynamic balance: the interference caused by integrating new knowledge is counteracted by the interference reduction achieved by ZiL, which is reflected in the fact that as the number of tasks increases, the norm of the RDB’s output doesn’t rise unboundedly but rather stabilizes after learning a certain number of tasks. This dynamic balance is crucial for maintaining the zero-shot generalizability of VLODMs.
@@ 4.1 Setup
Datasets. We conduct our experiments on the COCO [21] datasets and the “Object Detection in the Wild (ODinW)” [18] benchmark. ODinW is a more challenging benchmark designed to test model performance under real-world scenarios. It comprises numerous sub-datasets from various domains for evaluation, such as Thermal (to detect objects in heat map images) and Aquarium (to detect different marine animals). Following GLIP [19], we use ODinW-13 datasets, they are labeled as Ae (Aerial Maritime Drone), Aq (Aquarium), Co (Cottontail Rabbits), Eg (Egohands), Mu (Mushrooms), Pa (Packages), Pv (Pascal VOC), Pi (Pistols), Po (Pothole), Ra (Raccoon), Sh (Shellfish), Th (Thermal Dogs and People), Ve (Vehicles). The 13 sub-datasets of ODinW-13 are trained sequentially, one by one, and are tested after all sub-datasets have been trained.
Metrics. As COCO is a dataset within the realm of everyday life, we use zero-shot COCO performance as a metric of the model’s zero-shot generalization ability across general domains. Additionally, we provide the average performance on the downstream ODinW-13 benchmark, with performance calculations made only after the model has been sequentially trained on all sub-datasets. We denote the zero-shot mAP on COCO as “ZCOCO” and the average mAP on ODinW-13 downstream tasks as “Avg”.
Implementation Details. In this paper, unless otherwise specified, we use the Grounding DINO [22] with the Swin-Tiny [24] backbone as the default original VLODM in our experiments. This Grounding DINO is pre-trained on Objects365 [33], GoldG [19], and Cap4M [19]. We also use OV-DINO [35] with the Swin-Tiny backbone as an additional validation to demonstrate ZiRa’s ability to generalize effectively across different DETR-based architectures. This OV-DINO model is pre-trained on the Objects365 [33], GoldG [19], and Cap1M [35] datasets. The learning order of the ODinW-13 sub-datasets is randomly shuffled. All results are obtained after running experiments three times with different random seeds. The hyperparameter experiments are also conducted on the ODinW-13 dataset and COCO, and the results are shown in the appendix.
Our proposed method is implemented with PyTorch and trained on two Nvidia RTX 3090 GPUs. Each downstream task is trained for a total of two epochs with a batch size of 2. For Grounding DINO, we employ an initial learning rate of $10^{-3}$, which decays to 0.1 times the original value after the first epoch to ensure effective convergence. For OV-DINO, we employ an initial learning rate of $10^{-4}$, which also decays to 0.1 times the original value after the first epoch to ensure effective convergence. AdamW is used as the optimizer, and the weight decay is $10^{-4}$. At the beginning of IVLOD, we initialize the parameters of the total RDB as zero.
Table 1: IVLOD results on ODinW-13 and ZCOCO. All results are based on the same Grounding DINO. “Shots” means how many samples are used when adaptation training.
@@
Table 2: IVLOD results on ODinW-13 and ZCOCO. All results are based on the same OV-DINO.
Shots Methods ZCOCO Avg Ae Aq Co Eg Mu Pa Pv Pi Po Ra Sh Th Ve
Zero Original Model 50.22 26.64 15.69 19.37 11.79 40.67 1.23 59.26 50.78 12.19 2.44 33.81 8.50 43.68 46.86
Full TFA 44.98 47.56 22.27 35.36 63.58 56.68 19.84 64.26 68.07 50.01 25.99 56.31 33.70 56.96 65.20
Full CL-DETR 34.52 45.28 23.07 27.90 46.81 67.72 25.39 66.56 59.12 28.04 29.48 64.71 25.58 68.28 56.05
Full AT 36.80 44.33 22.46 27.42 45.78 67.10 24.10 65.08 58.17 26.65 28.88 64.46 24.60 67.30 54.27
Full iDETR 37.71 46.91 24.09 30.56 49.16 68.49 26.05 67.45 61.85 30.75 29.50 66.83 27.11 71.27 56.69
Full ZiRa 49.07 50.21 23.37 40.43 69.02 66.21 20.74 58.51 67.74 46.71 34.13 64.62 39.44 58.01 63.81
4.2 Comparison with Existing Methods
We first compare the above methods implemented on the Grounding DINO model. Experiments are conducted under both few-shot and full-shot settings. Since CL-DETR and OW-DETR are not designed for few-shot IOD, we only compare them under the full-shot setting. The results presented in Tab. 1 demonstrate that ZiRa consistently outperforms existing IOD methods in terms of “Avg” on downstream tasks. Moreover, ZiRa exhibits remarkable performance in preserving the zero-shot generalization ability of VLODMs under both the few-shot and full-shot settings. In particular, ZiRa surpasses CL-DETR and iDETR by substantial margins, with improvements of 13.91 and 8.74 AP under the full-shot setting in terms of “ZCOCO”, respectively.
We then compare ZiRa with existing methods implemented on the OV-DINO model. The results in Tab. 2 show that ZiRa outperforms existing IOD methods in terms of "Avg" on downstream tasks. ZiRa also demonstrates excellent performance in preserving the zero-shot generalization ability of VLODMs under the full-shot setting. Specifically, ZiRa surpasses CL-DETR and iDETR by significant margins, with improvements of 14.55 and 11.36 AP in the full-shot setting for "ZCOCO," respectively.
These results emphasize ZiRa’s effectiveness in maintaining the zero-shot AP and its potential as a superior solution for IOD. Distinctively, ZiRa, unlike CL-DETR and OW-DETR, requires minimal extra memory for a few branches without storing image exemplars and the model copy, making it more memory-efficient than other IOD methods.
4.3 Ablation Study
Table 3: Main ablation results.